# Perpetual Future Contracts in Centralized and Decentralized Exchanges: Mechanism and Traders’ Behavior

### v2

@@ 5.2 Empirical Approaches
Previous research has consistently shown that variables related to traders’ activities, such as trading volumes and open interests, exhibit significant correlation [33], suggesting their predictability based on the previous values. Consequently, these variables are often analyzed by dividing them into expected and unexpected components. The expected component, constructed from lagged variables, encompasses information about current trends, while the unexpected component captures unforeseen changes in traders’ behavior. This concept is illustrated in the following equation:
$\hat{\varepsilon}_{t}=\text{ Activity }_{t}-E\left\{\text{ Activity }_{t}\mid\text{ Activity }_{t-\tau},\tau=1,\ldots 10\right\},$ (28)
[…]
Beyond traditional trading activity variables like trading volumes and open interests, this study also assesses the conditional effects of leverage and liquidation on price volatility. We decompose these four types of variables into expected and unexpected components to differentiate the effects of hedgers and speculators, who vary in their trading willingness. Following the methodology of Bessembinder and Seguin (1993) [33], we partition trading volume, open interest, volume of liquidated positions, and daily average leverage using the Autoregressive Integrated Moving Average (ARIMA$(p,k,q)$) model. The ARIMA$(p,k,q)$ model, which predicts future values based on past data smoothed by lagged moving averages, allows us to designate the model’s fitted values as the expected component and its residuals as the unexpected component. For the ARIMA$(p,k,q)$ model, the existence of a unit root determines $k$, in which case a variable without a unit root is decomposed using ARIMA$(p,0,q)$, while a non-stationary variable is decomposed using ARIMA$(p,1,q)$  [^22] Readers can refer to [38] for details about ARIMA$(p,k,q)$ model.. The corresponding series is differenced $k$ times as an Autoregressive Moving Average (ARMA$(p,q)$) process. The values of $p$ and $q$ are selected based on the Akaike Information Criterion. The existence of unit roots in each time series is tested using the Augmented Dickey–Fuller (ADF) test, with results detailed in Appendix.
@@
$\hat{\sigma}_{t}=$ $\mu+\sum_{i=1}^{m}\phi_{i}\hat{\sigma}_{t-i}+\sum_{j=1}^{3}\alpha_{j}EA_{j,t}+\sum_{j=1}^{3}\beta_{j}UA_{j,t}+$ (29)
$\sum_{k=1}^{2}\gamma_{k}EL_{k,t}+\sum_{k=1}^{2}\lambda_{j}UL_{k,t}+\varepsilon_{t},$
@@
In addition to assessing how traders’ activities and liquidations respond to price volatility, we also examine speculation, as indicated by Eq. (25) and Eq. (26), using the following specification:
$\hat{\sigma}_{t}=$ $\mu+\sum_{i=1}^{m}\phi_{i}\hat{\sigma}_{t-i}+\sum_{j=1}^{3}\alpha_{j}TA_{j,t}+\beta_{SI}\mathcal{SI}_{t}+$ (30)
$\beta_{LIQS}\mathcal{LIQ}_{\text{short },t}+\beta_{LIQL}\mathcal{LIQ}_{\text{long },t}+\varepsilon_{t},$
where $\beta_{SI}$, $\beta_{LIQS}$, and $\beta_{LIQL}$ measure whether speculation is stronger in respond to the more volatile change in price. Among those, $\beta_{LIQS}$ and $\beta_{LIQL}$ indicate whether a short or long position is more likely to be liquidated than usual when the market become more volatile. $TA_{j,t}$ represent trading activities, i.e., $j$ represents trading volume, open interest on short positions, and open interest on long positions. Considering the considerable correlation between trading activities and price volatility, we include trading activities in Eq. (30) but do not decompose them as those are not the variables of interest.
Anticipating that traders are inclined to modify their leverage strategies in response to heightened liquidation risks under increased volatility of the underlying asset, this study incorporates an analysis of leverage adjustments. Research focusing on centralized exchanges (CEXs) often omits leverage considerations due to the general unavailability of such data [35, 21]. However, leveraging transaction-level data from GMX and GNS, this investigation calculates and integrates the daily average leverages for both long and short positions into the analytical framework. This is achieved by adapting the existing volatility estimation model, as delineated in Equation (29):
@@
In addition to exploring how average leverage changes with price volatility, we are also wondering how trading volume in different leverage level contribute to the change in average leverage level. To this end, we propose the following specification:
$\hat{\sigma}_{t}=$ $\mu+\sum_{i=1}^{m}\phi_{i}\hat{\sigma}_{t-i}+\sum_{j=1}^{2}\alpha_{j}EA_{j,t}+\sum_{j=1}^{2}\beta_{j}UA_{j,t}+\sum_{k=1}^{2}\gamma_{k}EL_{k,t}+$ (32)
$\sum_{j=1}^{2}\lambda_{j}UL_{k,t}+\sum_{k=1}^{10}\psi_{k}ELC_{k,t}+\sum_{j=k}^{10}\kappa_{j}ULC_{k,t}+\varepsilon_{t},$
In this equation, trading volumes are categorized into ten distinct groups based on their leverage levels: less than 1X, 1X to 5X, 5X to 10X, 10X to 15X, 15X to 20X, 20X to 25X, 25X to 30X, 30X to 35, 35X to 40X, 40X to 45X, 45X to 50X, and more than 50X. $ELC_{k,t}$ and $ULC_{k,t}$ denote the expected and unexpected portions of trading volume within each leverage category $k$ To avoid issues of multicollinearity, total trading volume is not included as a separate variable in $EA_{j,t}$ and $UA_{j,t}$, , given that the aggregated trading volume across all leverage categories equates to the total trading volume.
5.3 Empirical Results and Discussion
5.3.1 Descriptive statistics and dynamics of metrics.
Table 3 delineates the descriptive statistics for trading volumes, open interests, liquidations, and leverage of the five perpetual future platforms under examination, while Figure 8 graphically depicts these metrics on a weekly basis. Centralized exchanges (CEXs) have historically exhibited preeminence in trading volume and open interest, yet decentralized exchanges (DEXs) have demonstrated a notable ascension since 2023, as detailed in Appendix  [^24] Refer to Appendix for a visual representation of trading volume, open interest, and liquidation exclusive to DEXs.. An analysis of average daily liquidation volumes reveals a predilection for higher liquidations in long positions over short positions, with the exception of Perpetual Protocol V2. In the cases of GMX and GNS, the open interest for long positions surpasses that for short positions, and the leverages for both orientations (long & short) are relatively equivalent (15.57 & 15.80 for GMX and 72.33 & 75.28 for GNS), logically inferring a higher liquidation volume for long positions. With liquidation standardized, the close proximate values of $\mathcal{LIQ}_{short}$ and $\mathcal{LIQ}_{long}$ for GMX (0.016 and 0.016) and GNS (0.035 and 0.032) suggest a comparable probability of liquidation for both long and short positions. A symmetrical pattern in liquidation is anticipated between long and short positions with similiar leverage levels and volumes. This symmetry is observed for Perpetual Protocol V2 and Bybit, but not for Binance. A two-sample t-test indicates that the discrepancy between $\mathcal{LIQ}_{short}$ and $\mathcal{LIQ}_{long}$ is statistically significant exclusively for Binance (p-value = 0.001), whereas it remains statistically insignificant for the other platforms, namely GMX (p-value = 0.997), GNS (p-value = 0.813), Perpetual Protocol V2 (p-value = 0.883), and Bybit (p-value = 0.184). Despite the absence of leverage data for Binance, the order book matching mechanism posits an equivalent leverage distribution between sellers and buyers, yet empirical observations contradict this equilibrium. This discrepancy remains inconclusive pending further data.
Additionally, Figure 8 reveals an anomaly for Binance, where liquidation volumes for both long and short positions were concentrated during the bullish period in the first half of 2021, peaking in March and declining post-June, yet maintaining a consistently low level throughout the cryptocurrency market downturn in 2022. This trend may be attributable to Binance’s historical allowance of up to 125×125$\times$125 × leverage, which was reduced to a maximum of 20×20$\times$20 × on July 26, 2021252525For further details, refer to Binance’s announcement on July 27, 2021: https://www.binance.com/en/support/announcement/updates-on-rules-of-binance-futures-leverage-for-new-accounts-2021-07-27-d6457e23eb2e42f2b9c3ce44f46f9a6d. [39], significantly diminishing the likelihood of liquidation events.
Table 3: Descriptive Statistics of Trading Volumes, Open Interest, Liquidation, and Leverage of Perpetual Futures
Perpetual Contracts on Bitcoin
Statistics GMX GNS Perpetual Protocol V2 Binance Bybit
Mean Trading Volume (M) 40.46 7.62 2.65 411.55 92.18
OI on Short (M) 10.43 20.29 6.51 2235.61 831.39
OI on Long (M) 18.31 26.38 4.13 2235.61 831.39
Liquidation on Short (K) 62.68 197.30 4.53 11875.03 4527.90
Liquidation on Long (K) 124.09 248.76 3.54 17128.46 4753.27
Leverage on Short (K) 15.57 72.33 - - -
Leverage on Long (K) 15.80 75.28 - - -
$\mathcal{SI}$ 5.74 0.93 1.40 0.24 0.13
$\mathcal{LIQ}_{short}$ 0.016 0.035 0.002 0.009 0.012
$\mathcal{LIQ}_{long}$ 0.016 0.032 0.002 0.013 0.013
Std. Dev. Trading Volume (M) 50.82 12.40 2.75 249.56 84.57
OI on Short (M) 10.88 11.45 3.31 898.54 519.86
OI on Long (M) 19.37 16.87 1.36 898.54 519.86
Liquidation on Short (K) 459.08 1406.25 27.11 24231.71 7933.40
Liquidation on Long (K) 947.45 609.53 18.11 49440.90 7263.63
Leverage on Short (K) 6.89 14.32 - - -
Leverage on Long (K) 6.89 21.34 - - -
$\mathcal{SI}$ 14.71 2.28 5.51 0.19 0.10
$\mathcal{LIQ}_{short}$ 0.173 0.359 0.017 0.022 0.024
$\mathcal{LIQ}_{long}$ 0.173 0.101 0.029 0.040 0.026
No. of observations 747 628 668 1147 1147
Note. This table reports the daily mean and standard deviation of trading volumes (in million USD), open interest (O.I. in million USD), short liquidations (in thousand USD), long liquidations (in thousand USD), $\mathcal{SI}$, $\mathcal{LIQ}_{\text{short }}$, $\mathcal{LIQ}_{\text{long }}$ of USD or USDT bitcoin perpetuals across GMX, GNS, Perpetual Protocol V2, Binance, and Bybit. Daily mean and standard deviation of average leverage on short and long positions (in thansand USD) are reported for perpetuals on GMX and GNS. The measure of daily average leverage is calculated by dividing the sum of daily transaction volume by the total amount of collateral supporting those transactions, which is the weighted average leverage of transactions in a day. We also conduct ADF test for each of the series, the results of which can be found in Appendix.
Figure 8: Trading Volumes, Open Interest, and Liquidated Volumes of Perpetual Futures on all the Five DEXs and CEXs. As CEXs dominate the market, we also create the figure solely illustrating DEXs, which is in Appendix.
Upon examination of the mean values for $\mathcal{SI}$ and $\mathcal{LIQ}$ delineated in Table 3 across various trading platforms, it is discernible that GMX exhibits the highest mean value for $\mathcal{SI}$ at 5.74, while GNS registers the maximal values for both $\mathcal{LIQ}_{short}$ (0.035) and $\mathcal{LIQ}_{long}$ (0.032) respectively. This suggests a relatively heightened level of speculative trading activity on GMX and GNS compared to other exchanges. Such a trend aligns with the operational parameters of GMX and GNS, which afford traders the opportunity to engage with higher leverage options and impose a lower threshold for maintenance margins relative to their counterparts. Specifically, GMX and GNS extend maximum leverage ratios of 100×100$\times$100 × and 150×150$\times$150 × respectively, in contrast to centralized exchanges (CEXs) like Binance and Bybit, which implement a tiered leverage system that inversely correlates the maximum allowable leverage with the size of the position  [^26] For maximum leverages allowed in each tier of position size on Binance and Bybit, readers can refer to https://www.binance.com/en-IN/futures/trading-rules/perpetual/leverage-margin and https://www.bybit.com/en-US/help-center/article/Risk-Limit-USDC-Contract. Since Perpetual Protocol V2 adopts cross-margining, the leverage for individual position is undefined.. The maintenance margin required by GMX is 1%, and the liquidation price on GNS is calculated by:
@@ 5.3 Empirical Results and Discussion
[…]
Effectively, traders on GNS trade with the lowest maintenance margin ratio among the five exchanges, where Perpetual Protocol V2 adopts the maintenance margin ratio of 6.25% and 0.4% for both Binance and Bybit in the tier of the smallest position  [^27] With Eq. (33), if a trader opens a long position with collateral of 50 USD and open price of 20,000 USD, he/she would be liquidated when the price drops to 19,824 USD, liquidated with a maintenance margin ratio of 0.1% equivalently. .
Figure 9: Trading Volume in Leverage Categories and Average Leverage of Daily Transactions (GMX). The upper panel shows the dynamics of the daily trading volume in each of the 12 leverage categories defined in Section 5.2.2, while the lower panel shows the daily average leverage for long and short positions.
Figure 10: Trading Volume in Leverage Categories and Average Leverage of Daily Transactions (GNS).The upper panel shows the dynamics of the daily trading volume in each of the 11 leverage categories defined in Section 5.2.2, i.e., there does not exist transaction with less than 1$\times$ leverage in our dataset on GNS, while the lower panel shows the daily average leverage for long and short positions.
Fig. 9 and Fig. 10 illustrate the dynamics of the daily trading volume in different leverage categories defined in Section 5.2.2 as well as the daily average leverage level for long and short positions respectively. GMX announced to rise its maximum leverage on 8 November 2021 [40], followed with significant trading volumes in leverages more than 35$\times$ after November 2021, as shown in the upper panel of Fig. 9. It shows that although trading volumes with different leverages vary across time, the level of daily average leverage oscillates within a narrow range. The two-sample t-test indicates the insignificant difference of leverage between long and short positions for GMX, with p-value of 0.515, while the difference is significant for GNS, with p-value of 0.004, in which case long positions are with higher leverages than short positions averagely for GNS. This finding is consistent with the record by BitMEX CEO Arthur Hayes [41], in which the average effective leverage for long positions is on average higher than that for short positions. Around April 2023, there was a sudden drop in average leverage on long positions for GNS, shown in Fig. 10. This change could be attributed to the huge drop in the daily value of collateral used for transactions on long positions during this period, as the average leverage is calculated with dividing the aggregated value of collateral by the aggregated transaction value each day.
Figure 11: Volume-normalized Liquidations for DEXs. The middle and bottom panel depict the fraction of liquidations on long and short positions over total trading volume across time respectively.
Fig. 11 shows the fraction of liquidations over total trading volume across time for DEXs, which is the liquidations standardized by trading volumes and also a proxy for evaluating the risk of an instrument [21]. As expected, GNS acts as the most risky DEX across time, for its liquidations standardized by trading volumes for both long and short positions consistently higher than that for GMX and Perpetual Protocol V2 overtime. The standardized liquidation of long (short) positions increases even when the Bitcoin price trended up (down), however, short after July 2022, a counter-intuitive pattern also noted by [21]. [21] mentions the increased volatility accompanied by the uptrend of Bitcoin price, which presents as high intra-day price volatility due to price expansion, as a potential explanation but without empirical evidence. In the rest of this section, we systematically examine the association between liquidation the price volatility of Bitcoin, explaining the pattern of liquidation as well as the asymmetry between long and short positions.
5.3.2 Empirical results and analysis for CEXs.
Table 4 aggregates regression results on Eq. (29) and Eq. (30). Coefficient estimates for expected and unexpected trading volume are positive across all the five exchanges, and are statistically significant except the one for expected trading volume on GNS. Estimated coefficients for the unexpected volume (the shocks) are uniformly higher than those for the unexpected volume, suggesting that one unit change in the unexpected portion of trading volumes has higher effect on the price volatility of Bitcoin than one unit change in the expected volume, consistent with the findings in the previous studies on traditional futures market [33, 34].
For CEXs, i.e., Binance and Bybit, all the estimated coefficients measuring the marginal effects of expected and unexpected open interest are negative and significant, while the estimated coefficients for the unexpected portion are also larger in magnitude than those for the expected portion. Since in CEXs each seller is matched with a buyer of the perpetual futures, aggregated value of long positions equals to that of short positions, so do the open interests. We only include open interests on short in Eq. (29), as open interests on long are of the same value. Since CEXs adopt the mechanism with order books, it relates to the determinants of market depth, which [1] defines as the volume of unanticipated order flow required to move price by one unit. When capital inflows into the market, the liquidity increases, which is reflected by higher level of open interests, making the price less volatile in response to new orders. Consequently, it is expected that when open interest is large the price volatility conditional on contemporaneous trading volume, which proxies market depth, would be lower. Therefore, those negative estimated coefficients imply that an increase in unexpected open interest, e.g., inflow of speculative capital, lessens the impact of trading volume shock (unexpected volume) on price volatility. In our case, for Binance, when a trader increase his/her position by 1 USD, the marginal effect of the 1 USD increase in trading volume on price volatility is $5.943\times 10^{-8}$, while mitigated by the effect of 1 USD increase in open interest ($-2.365\times 10^{-11}$), resulting in a smaller aggregated effect ($5.943\times 10^{-8}-2.365\times 10^{-11}$). On the one hand, estimated coefficients for trading volumes and open interests have opposite signs, while on the other hand, whether the marginal effect of volume on price volatility is enlarged or mitigated depends on whether open interest is reduced or increased. Trades closing, decreasing, or liquidating positions actually reduce open interest, enlarging the effect of trading volume on price volatility  [^28] Trades opening positions or rising positions increases open interests.. Price would move with larger distance, as less liquidity reacts to absorb a trade. This is reflected by that seven out of eight coefficients for liquidated volume in Binance and Bybit are significant and positive, with expected liquidated volume on short positions for Bybit as the exception. Thus, ceteris paribus, an unexpected rise in liquidation reduces market depth, via lowering open interest and liquidity. Considering the abnormality in the dynamics of liquidation data for Binance, we replicate the regressions in Eq. (29) and Eq. (30) for two periods dichotomized by 1 May 2021, the date around which liquidations on Binance suddenly dropped and became consistently low. The estimating results and analysis for the two periods separately are included in Appendix. Traders’ behavior reflected in the two periods is almost the same with that reflected in Table 4.
Table 4: Regression Results on Eq. (29) and Eq. (30)
Perpetual Contracts on Bitcoin
Variables GMX GNS Perpetual Protocol V2 Binance Bybit
Panel a: Regression Results on Eq. (29):
Intercept 1.438e-12 (16.82)*** 0.017 (8.82)*** 0.009 (3.48)*** 0.011 (6.27)*** 2.145 (0.15)
Lagged volatility 0.256 (8.10)*** 0.371 (10.57)*** 0.252 (8.19)*** 0.200 (10.16)*** 0.210 (17.19)***
@@
Expected trading volume 1.186e-10 (6.30)*** 1.222e-11 (0.19) 2.59e-09 (8.76)*** 3.06e-08 (12.87)*** 6.041e-08 (8.91)***
Unexpected trading volume 1.29e-10 (9.83)*** 4.53e-10 (6.65)*** 3.62e-09 (14.42)*** 5.943e-08 (24.07)*** 1.409e-07 (15.62)***
Expected OI on short -4.765e-10 (-7.36)*** -2.059e-10 (-1.99)** -1.678e-09 (-4.40)*** -1.11e-12 (-2.24)*** -8.55e-12 (-8.69)***
Unexpected OI on short -2.808e-10 (-3.27)*** -4.654e-10 (-3.08)*** 1.125e-08 (4.53)*** -2.365e-11 (-8.93)*** -3.138e-11 (-4.65)***
Expected OI on long -1.88e-10 (-5.99)*** 2.98e-11 (0.44) 3.569e-09 (3.98)*** - -
Unexpected OI on long -1.75e-10 (-2.16)** -4.261e-11 (-0.29) 1.481e-09 (0.161) - -
@@
Expected liquidated volume on short positions 2.449e-09 (1.09) -1.483e-09 (-1.26) -5.983e-07 (-2.09)** 2.864e-10 (5.46)*** -2.832e-10 (-2.32)**
Unexpected liquidated volume on short positions 2.007e-09 (1.61) 3.386e-10 (0.835) -3.032e-08 (-1.61) 1.655e-10 (6.95)*** 5.098e-10 (7.22)***
Expected liquidated volume on long positions 1.82e-07 (17.71)*** 7.295e-09 (3.26)*** 5.008e-07 (4.16)*** 8.534e-11 (3.00)*** 1.672e-09 (11.96)**
Unexpected liquidated volume on long positions 2.593e-09 (5.08)*** 1.448e-09 (1.46) 1.482e-07 (5.39)*** 7.143e-11 (6.82)*** 1.278e-09 (16.62)***
Adjusted $R^{2}$ 0.375 0.310 0.507 0.647 0.649
AIC -4381 -3670 -4067 -6580 -6664
No. of obs. 747 628 668 1147 1147
Panel b: Regression Results on Eq. (30):
Intercept 0.021 (18.02)*** 0.017 (9.88)*** 0.001 (0.51) -0.003 (-1.46) 0.015 (10.00)***
Lagged volatility 0.260 (8.40)*** 0.382 (10.69)*** 0.256 (8.18)*** 0.224 (11.12)*** 0.189 (8.95)***
Trading Activity:
Trading volume 1.368e-10 (10.58)*** 3.007e-10 (4.72)*** 3.335e-09 (15.79)*** 3.66e-08 (13.28)*** 1.096e-07 (9.17)***
OI on short -4.31e-10 (-8.66)*** -2.802e-10 (-3.33)*** -1.098e-09 (-2.89)*** 2e-12 (2.38)** -7.609e-12 (-5.60)***
OI on long -1.966e-10 (-6.22)*** 6.244e-11 (1.10) 3.74e-09 (3.95)*** - -
$\mathcal{SI}$ and $\mathcal{LIQ}$:
$\mathcal{SI}$2.597e-05 (-0.63) -0.001 (-1.79)* 3.000e-04 (3.19)*** 0.020 (4.07)*** 0.004 (0.40)
$\mathcal{LIQ}_{\text{short }}$ 0.004 (1.48) -0.001 (-0.816) -0.065 (-2.29)** 0.184 (7.30)*** 0.169 (7.58)***
$\mathcal{LIQ}_{\text{long }}$ 0.015 (5.78)*** 0.017 (2.75)*** 0.019 (1.19) 0.156 (11.73)*** 0.34 (0.40)
Adjusted $R^{2}$ 0.380 0.282 0.466 0.600 0.600
AIC -4389 -3649 -4018 -6435 -6511
No. of obs. 747 628 668 1147 1147
Notes: Panel a of this table reports the regression results of Eq. (29) while Panel b reports the regression results of Eq. (30). For perpetual futures on both DEXs and CEXs, contract size is undefined and traders trade perpetual futures with arbitrary units of Bitcoins, so trading volumes and open interests are measured in USD. Since for DEXs each future contract is matched with a buyer and a seller, resulting in the same dollar value of open interests for long and short, we regress price volatility only on open interest on short. This is only for notation while does not mean CEXs are only with short positions, because the dollar value of open interest on short is always equals to that on long. In each cell, the t-statistics are in the parentheses. ***, ***** *, and ******* * * denote significance at 0.1, 0.05, and 0.01 level, respectively.
5.3.3 Empirical results and analysis for Perpetual Protocol V2 (with VAMM Model).
@@
In CEXs, traders are matched by the market maker and provide liquidity. Different from the case of CEXs, traders in DEXs are matched with liquidity pools, in which liquidity providers act passively as the counterparty. Traders’ open interest thus does not necessarily influence the market depth in the same way as in CEXs and the theory of market depth by [1] for traditional futures market is inapplicable in this context. For Perpetual Protocol V2, which adopts VAMM as the pricing mechanism, liquidity providers deposit stable coins in the vault based on which the clearing house mints virtual tokens and places those virtual tokens into the Uniswap v3 AMM  [^29] Details of how VAMM works in Perpetual Protocol V2 can be found in https://support.perp.com/hc/en-us/articles/9594157347481-How-It-Works.. For illustration, assume the price of the Bitcoin perpetual is determined by the exchange function defined in Section 3.8:
@@
The implication of Uniswap v3 AMM on pricing volatility is consistent with our empirical findings. In Table 4, the coefficient for expected open interest on short is estimated to be significantly negative ($-1.678\times 10^{-9}$) while that on long is estimated to be significantly positive ($3.569\times 10^{-9}$), indicating a negative association between expected open interest on short and price volatility while a positive association for long. Since the expected portion of open interest reflects traders’ long-run behavior, e.g., hedgers, in cope with the trend, the change could be mainly be described by Eq. (34). When the expected open interest on short rises, the clearing house retrieve vBTC from traders’ account and credit to the balance of the liquidity pool. The increased vBTC balance in the liquidity pool helps decrease the rate of pricing change and thus reducing volatility according to Eq. (36). Besides, when expected open interest on long increases, the clearing house withdraw vBTC from the pool and credit to traders’, decreasing vBTC balance in the liquidity pool and thus increasing price volatility.

### v4

@@ 2 Data and Methodology
Previous research has consistently shown that variables related to traders’ activities, such as trading volumes and open interests, exhibit significant correlation [5], suggesting their predictability based on the previous values. Consequently, these variables are often analyzed by dividing them into expected and unexpected components. The expected component encompasses information about current trends, while the unexpected component captures unforeseen changes in traders’ behavior. Hedgers, who adjust positions infrequently with a long-term focus, exhibit predictable behavior captured in the expected component. In contrast, speculators, who frequently adjust positions to manage risk, are represented by the unpredictable, unexpected component of activity. Following the methodology of Bessembinder and Seguin (1993) [5], we partition trading volume, open interest, volume of liquidated positions, and daily average leverage using ARIMA$(p,k,q)$ model, where fitted values are regarded as the expected component and its residuals as the unexpected component.
@@
$\sum_{k=1}^{2}\lambda_{j}UL_{k,t}+\varepsilon_{t},$
@@
Anticipating that traders are inclined to modify their leverage strategies in response to heightened liquidation risks under increased price volatility of the underlying asset, this study incorporates an analysis of leverage adjustments. Research focusing on centralized exchanges (CEXs) often omits leverage considerations due to the general unavailability of such data [22, 20]. However, leveraging transaction-level data from GMX and GNS, this investigation calculates and integrates the daily average leverages for both long and short positions into the analytical framework. This is achieved by adapting the existing volatility estimation model, as delineated in Equation (2):
@@
3 Empirical Results
3.1 Empirical results and analysis for LOB Model (Binance and Bybit)
Table 1: Regression Results on Eq. (2)
@@ 3.1 Empirical results and analysis for LOB Model (Binance and Bybit)
Binance Bybit [^8] Since for CEXs each future contract is matched with a buyer and a seller, resulting in the same dollar value of open interests for long and short, we regress price volatility only on open interest on short. Perpetual Protocol V2 GMX GNS
Intercept [^9] In each cell, the t-statistics are in the parentheses. *, **∗ ∗, and ***∗ ∗ ∗ denote significance at 0.1, 0.05, and 0.01 level, respectively. 0.011 (6.27)*** 2.145 (0.15) 0.009 (3.48)*** 1.438e-12 (16.82)*** 0.017 (8.82)***
Lagged volatility 0.200 (10.16)*** 0.210 (17.19)*** 0.252 (8.19)*** 0.256 (8.10)*** 0.371 (10.57)***
@@
Expected trading volume 3.06e-08 (12.87)*** 6.041e-08 (8.91)*** 2.59e-09 (8.76)*** 1.186e-10 (6.30)*** 1.222e-11 (0.19)
Unexpected trading volume 5.943e-08 (24.07)*** 1.409e-07 (15.62)*** 3.62e-09 (14.42)*** 1.29e-10 (9.83)*** 4.53e-10 (6.65)***
Expected OI on short -1.11e-12 (-2.24)*** -8.55e-12 (-8.69)*** -1.678e-09 (-4.40)*** -4.765e-10 (-7.36)*** -2.059e-10 (-1.99)**
Unexpected OI on short -2.365e-11 (-8.93)*** -3.138e-11 (-4.65)*** 1.125e-08 (4.53)*** -2.808e-10 (-3.27)*** -4.654e-10 (-3.08)***
Expected OI on long - - 3.569e-09 (3.98)*** -1.88e-10 (-5.99)*** 2.98e-11 (0.44)
Unexpected OI on long - - 1.481e-09 (0.161) -1.75e-10 (-2.16)** -4.261e-11 (-0.29)
@@
Expected liquidated volume on short positions 2.864e-10 (5.46)*** -2.832e-10 (-2.32)** -5.983e-07 (-2.09)** 2.449e-09 (1.09) -1.483e-09 (-1.26)
Unexpected liquidated volume on short positions 1.655e-10 (6.95)*** 5.098e-10 (7.22)*** -3.032e-08 (-1.61) 2.007e-09 (1.61) 3.386e-10 (0.835)
Expected liquidated volume on long positions 8.534e-11 (3.00)*** 1.672e-09 (11.96)** 5.008e-07 (4.16)*** 1.82e-07 (17.71)*** 7.295e-09 (3.26)***
Unexpected liquidated volume on long positions 7.143e-11 (6.82)*** 1.278e-09 (16.62)*** 1.482e-07 (5.39)*** 2.593e-09 (5.08)*** 1.448e-09 (1.46)
Adjusted $R^{2}$ 0.600 0.600 0.466 0.380 0.282
AIC -6435 -6511 -4018 -4389 -3649
No. of obs. 1147 1147 668 747 628
table 1 aggregates regression results on Eq. (2). Coefficient estimates for expected and unexpected trading volume are positive across all the five exchanges, consistent with the findings in the previous studies on traditional futures market [5, 24].
For exchanges of LOB Model, i.e., Binance and Bybit, all the estimated coefficients measuring the marginal effects of expected and unexpected open interest are negative and significant, while the estimated coefficients for the unexpected portion are also larger in magnitude than those for the expected portion. Since in these exchanges each seller is matched with a buyer of the perpetual futures, aggregated value of long positions equals to that of short positions, so do the open interests. We only include open interests on short in Eq. (2), as open interests on long are of the same value. Since exchanges of LOB Model adopt the limit-order-books, it relates to the determinants of market depth, which Kyle (1985)[13] defines as the volume of unanticipated order flow required to move price by one unit. When capital inflows into the market, the liquidity increases, which is reflected by higher level of open interests, making the price less volatile in response to new orders. Consequently, it is expected that when open interest is large the price volatility conditional on contemporaneous trading volume, which proxies market depth, would be lower. Therefore, those negative estimated coefficients imply that an increase in unexpected open interest, e.g., inflow of speculative capital, lessens the impact of trading volume shock (unexpected volume) on price volatility. In our case, for Binance, when a trader increase his/her position by 1 USD, the marginal effect of the 1 USD increase in trading volume on price volatility is $5.943\times 10^{-8}$, while mitigated by the effect of 1 USD increase in open interest ($-2.365\times 10^{-11}$), resulting in a smaller aggregated effect ($5.943\times 10^{-8}-2.365\times 10^{-11}$). On the one hand, estimated coefficients for trading volumes and open interests have opposite signs, while on the other hand, whether the marginal effect of volume on price volatility is enlarged or mitigated depends on whether open interest is reduced or increased. Trades closing, decreasing, or liquidating positions actually reduce open interest, enlarging the effect of trading volume on price volatility  [^10] Trades opening positions or rising positions increases open interests.. Price would move with larger distance, as less liquidity reacts to absorb a trade. This is reflected by that seven out of eight coefficients for liquidated volume in Binance and Bybit are significant and positive, with expected liquidated volume on short positions for Bybit as the exception. Thus, all else being equal, an unexpected rise in liquidation reduces market depth, via lowering open interest and liquidity.
3.2 Empirical results and analysis for VAMM Model (Perpetual Protocol V2).
@@
Different from the case of LOB Model, where traders are matched by the market maker and provide liquidity, traders in exchanges of VAMM Model are matched with liquidity pools, in which liquidity providers act passively as the counterparty. Traders’ open interest thus does not necessarily influence the market depth in the same way as in LOB Model and the theory of market depth by Kyle (1985)[13] for traditional futures market is inapplicable in this context. For Perpetual Protocol V2, which adopts VAMM as the pricing mechanism, liquidity providers deposit stable coins in the vault based on which the clearing house mints virtual tokens and places those virtual tokens into the Uniswap v3 AMM  [^11] Details of how VAMM works in Perpetual Protocol V2 can be found in https://support.perp.com/hc/en-us/articles/9594157347481-How-It-Works.. For illustration, assume the price of the Bitcoin perpetual is determined by the exchange function:
@@
The implication of Uniswap v3 AMM on pricing volatility is consistent with our empirical findings. In Table 1, the coefficient for expected open interest on short is estimated to be significantly negative ($-1.678\times 10^{-9}$) while that on long is estimated to be significantly positive ($3.569\times 10^{-9}$), indicating a negative association between expected open interest on short and price volatility while a positive association for long. Since the expected portion of open interest reflects traders’ long-run behavior, e.g., hedgers, in cope with the trend, the change could be mainly be described by Eq. (4). When the expected open interest on short rises, the clearing house retrieve vBTC from traders’ account and credit to the balance of the liquidity pool. The increased vBTC balance in the liquidity pool helps decrease the rate of pricing change and thus reducing price volatility according to Eq. (6). Besides, when expected open interest on long increases, the clearing house withdraw vBTC from the pool and credit to traders’, decreasing vBTC balance in the liquidity pool and thus increasing price volatility.